# Learning Darkness Clue Prompting for Nighttime UAV Tracking

### v1

@@ Abstract
Existing nighttime unmanned aerial vehicle (UAV) trackers follow an “Enhance-then-Track” architecture - first using a light enhancer to brighten the nighttime video, then employing a daytime tracker to locate the object. This separate enhancement and tracking fails to build an end-to-end trainable vision system. To address this, we propose a novel architecture called Darkness Clue-Prompted Tracking (DCPT) that achieves robust UAV tracking at night by efficiently learning to generate darkness clue prompts. Without a separate enhancer, DCPT directly encodes anti-dark capabilities into prompts using a darkness clue prompter (DCP). Specifically, DCP iteratively learns emphasizing and undermining projections for darkness clues. It then injects these learned visual prompts into a daytime tracker with fixed parameters across transformer layers. Moreover, a gated feature aggregation mechanism enables adaptive fusion between prompts and between prompts and the base model. Extensive experiments show state-of-the-art performance for DCPT on multiple dark scenario benchmarks. The unified end-to-end learning of enhancement and tracking in DCPT enables a more trainable system. The darkness clue prompting efficiently injects anti-dark knowledge without extra modules. Code and models will be released.
@@ I Introduction
As a fundamental vision capability for unmanned aerial robots, UAV-based object tracking has attracted extensive research in academia and industry, since it can facilitate a variety of downstream intelligent applications, e.g., traffic monitoring [1], aerial cinematography [2], and search and rescue mission [3]. Recent advances [4, 5, 6] in visual object tracking achieve promising results with increasingly deep neural networks [7, 8, 9] and large-scale tracking datasets [10, 11, 12]. However, existing top-notch trackers are mainly served for daytime scenarios. When faced with more challenging light conditions (e.g., the night falls), these methods often suffer from severe performance degradation or even fail to work. This is because the discriminative clues (e.g., color and geometric information) are weakened and more noise speckles [13] are brought by onboard cameras in nighttime circumstances, the tracker is unable to extract valid representational features for accurate localization. Therefore, exploring effective nighttime tracking methods will promote the generalizability and survivability of UAV vision systems.
@@
Several studies have been conducted to empower the UAV system with the ability to “see” in the dark. Fu et al. [14] propose a light enhancer named HighlightNet to brighten up the potential object region for UAV trackers. This “Enhance-then-Track” paradigm (Fig. 1 (a)) is also adopted in [15, 16]. They mainly focus on designing an effective light enhancer and employ off-the-shelf daytime trackers to output the final tracking results. On the other hand, Ye et al. [17] introduce domain adaptation (Fig. 1 (b)) for nighttime UAV tracking. They generate nighttime training samples and adversarially train a model for narrowing the gap between day and night circumstances. Despite gaining improvements, current solutions for nighttime tracking still have significant limitations. i) Executing image enhancement before tracking makes UAV vision system over-reliance on the extra trained enhancer, and separating the nighttime tracking into two sub-processes masks is not conducive to building an end-to-end trainable architecture. ii) Domain adaptation requires abundant data for training and high-quality target domain samples are scarce for nighttime domain learning. iii) The intrinsic relationship between daytime tracker and nighttime tracker is overlooked, and the potential for employing daytime tracker in nighttime scenarios is not well exploited.
@@
Recently, prompt learning has attracted much attention, extending from the Natural Language Processing (NLP) to the vision tasks [18, 19, 20]. Typically, the foundation model is frozen, and only a few tunable parameters are added for learning valid prompts. This approach demonstrates promising results and efficiencies. Drawing inspiration from these works, we formulate nighttime UAV tracking as a prompt learning problem where the goal is to mine valid darkness clue prompts for a well-trained daytime tracker, so that the parameter and computational cost are constrained and the nighttime performance is maximized. In this work, we propose a novel nighttime UAV tracker (termed DCPT) which dwells on learning darkness clue prompting for low-light circumstances (Fig. 1 (c)). Specifically, to effectively facilitate the discovery and mining of clues in darkness, we design a darkness clues prompter (DCP) by introducing the back-projection structure which shows favorable performance in image restoration (e.g., image super-resolution [21]). DCP propagates the darkness clue prompts across all semantic layers of the foundation tracker. Moreover, we design a gated feature aggregation (GFA) mechanism to efficiently fuse bottom-up prompts and enable the complementary integration of learned prompts and information from the foundation model. Ultimately, DCPT can effectively inject learned darkness clue prompts into a frozen daytime model with only a small number of prompt learning-related parameters, obtaining superior nighttime tracking performance. We summarize our major contributions as follows:
@@
Extensive experiments on four nighttime tracking benchmarks validated the effectiveness of DCPT. Qualitative and quantitative results demonstrate the superiority of DCPT as a nighttime tracker for aerial robots.
@@ II-A Nighttime UAV Tracking
The emergence of deep-learning technologies and efforts by researchers push visual object tracking to new frontiers[22, 23, 5, 6]. Impressive tracking performance was also achieved on the UAV platforms [24, 4]. However, these trackers, which are mainly designed for daytime scenarios, often suffer severe performance degradation or even fail to work when faced with common but challenging nighttime scenarios. This is because the nighttime circumstances suffer from loss of detailed information, accompanying noise and low contrast and brightness. Therefore, nighttime UAV tracking has attracted increasing attention. A straightforward manner is to complete nighttime tracking by an ”Enhance-then-Track” process. Specifically, researchers design low light enhancers [25, 15, 14] for nighttime scenarios, and use existing daytime UAV trackers to tracking in the enhanced sequences. Li et al. [16] propose to integrate a low-light image enhancer into a CF-based tracker for robust tracking at night. Similarly, DarkLighter [15] and HighlightNet [14] also design low-light image enhancers to alleviate the influence of extreme illumination and highlight the potential objects, respectively. Although effective, the main drawbacks of such approaches are that the nighttime tracker requires an additional trained light enhancer for preprocessing, incurring extra computational costs, and the separation of enhancer and tracker is not conducive to building an end-to-end trainable nighttime UAV tracking method. Another approach is to adopt domain adaptation to transfer daytime tracker to nighttime scenarios. To obtain the tracking capabilities in the dark, UDAT [17] proposes to align image features from daytime and nighttime domains by the transformer-based bridging layer, in this way, the tracking capabilities on the daytime domain are somewhat transferred to the nighttime domain. Unfortunately, this approach requires more training costs and the lack of high-quality target domain data for tracking also limits its enhancement.
@@ III-A Overview
The overall architecture is shown in Fig. 2. To summarize, DCPT injects the learned darkness clue prompts into the daytime tracker and stimulates its tracking potential in low-light circumstances. These learned darkness clue prompts are propagated from preceding prompts and foundation feature flows. Specifically, the gated feature aggregation (GFA) mechanism is designed for controlling the weight of these different information sources. Intuitively, we believe that the daytime tracker has excellent tracking capabilities on generic scenarios while just lacking specific adaptations for nighttime scenarios. Therefore, we only tune parameters that are related to darkness clue prompt generation instead of fine-tuning the entire model. The DCPT paradigm maximizes the capabilities inherited from daytime tracker trained on large-scale datasets [11, 27, 12], avoiding overfitting on limited nighttime tracking data but gaining nighttime-specific tracking capabilities through darkness clue prompt learning.
III-B Daytime Base Tracker
High-performance base tracker is the guarantee of the success of a nighttime tracker, and previous nighttime tracking methods [16, 14, 15] all chose cutting-edge trackers [28, 4, 23] as their daytime base trackers. Hereby we construct a pure transformer daytime tracker as our strong foundation model. We make it as simple as possible to clearly show the effectiveness of our darkness clue prompting tracking paradigm. Specifically, the base tracker is only composed of two parts: a backbone and a prediction head. The backbone consists of stacked vision transformer (ViT) layers and propagating the inputs in a one-stream style like the method in [6]. First, both template and search region are embedded and flattened to 1D tokens with added positional embedding :

### v4

# DCPT: Darkness Clue-Prompted Tracking in Nighttime UAVs

@@ Abstract
Existing nighttime unmanned aerial vehicle (UAV) trackers follow an “Enhance-then-Track” architecture - first using a light enhancer to brighten the nighttime video, then employing a daytime tracker to locate the object. This separate enhancement and tracking fails to build an end-to-end trainable vision system. To address this, we propose a novel architecture called Darkness Clue-Prompted Tracking (DCPT) that achieves robust UAV tracking at night by efficiently learning to generate darkness clue prompts. Without a separate enhancer, DCPT directly encodes anti-dark capabilities into prompts using a darkness clue prompter (DCP). Specifically, DCP iteratively learns emphasizing and undermining projections for darkness clues. It then injects these learned visual prompts into a daytime tracker with fixed parameters across transformer layers. Moreover, a gated feature aggregation mechanism enables adaptive fusion between prompts and between prompts and the base model. Extensive experiments show state-of-the-art performance for DCPT on multiple dark scenario benchmarks. The unified end-to-end learning of enhancement and tracking in DCPT enables a more trainable system. The darkness clue prompting efficiently injects anti-dark knowledge without extra modules. Code is available at https://github.com/bearyi26/DCPT.
@@ I Introduction
Visual object tracking from unmanned aerial vehicles (UAVs) is an essential capability of aerial robotic vision, enabling various downstream applications such as traffic monitoring [1], aerial cinematography [2], and search and rescue [3]. While recent advances using deep neural networks [7, 8, 9] and large-scale datasets [10, 11, 12] have achieved promising tracking performance in daytime conditions, state-of-the-art trackers [4, 5, 6] still struggle in more challenging nighttime environments. When faced with more challenging light conditions (e.g., the night falls), these approaches often suffer from severe performance degradation or even fail to work. This is mainly because discriminative visual cues like color and geometry are diminished at night, and onboard cameras introduce more noise and image degradation under low illumination. As a result, existing trackers fail to extract robust features for accurate target localization at night. Therefore, to fully realize the potential of UAV vision, it is imperative to explore effective night-time tracking techniques, which will promote the versatility and survivability of UAV vision systems.
@@
Several studies have been conducted to equip UAV systems with the capacity to "see" in low-light conditions. For example, Fu et al. [14] introduced a light enhancer called "HighlightNet" designed to illuminate specific target areas for UAV trackers. This “Enhance-then-Track" paradigm (Fig. 1 (a)) is also adopted in other studies [15, 16]. They mainly focus on designing a light enhancer and employ off-the-shelf daytime trackers to output the final tracking results. On the other hand, Ye et al. [17] introduce domain adaptation (Fig. 1 (b)) for nighttime UAV tracking. They generate nighttime training samples and adversarially train a model for narrowing the gap between day and night circumstances. Despite gaining improvements, current solutions for nighttime tracking still have significant limitations. i) Performing image enhancement before tracking makes UAV vision system over-reliance on the extra trained enhancer, and separating the nighttime tracking into two sub-processes masks is not conducive to building an end-to-end trainable architecture. ii) Domain adaptation requires abundant data for training and high-quality target domain samples are scarce for nighttime domain learning. iii) The intrinsic relationship between daytime tracker and nighttime tracker is overlooked, and the potential for employing daytime tracker in nighttime scenarios is not well exploited.
@@
Recently, prompt learning has attracted much attention, extending from Natural Language Processing (NLP) to the vision tasks [18, 19, 20]. Typically, the foundation model is frozen, and only a few tunable parameters are added for learning valid prompts. This approach demonstrates promising results and efficiencies. Drawing inspiration from these works, we formulate nighttime UAV tracking as a prompt learning problem where the goal is to mine valid darkness clue prompts for a well-trained daytime tracker, so that the parameter and computational cost are constrained and the nighttime performance is maximized. In this work, we propose a novel nighttime UAV tracker (termed DCPT) that dwells on learning darkness clue prompting for low-light circumstances (Fig. 1 (c)). Specifically, to effectively facilitate the discovery and mining of clues in darkness, we design a darkness clues prompter (DCP, in III-C) by introducing the back-projection structure which shows favorable performance in image restoration (e.g., image super-resolution [21]). DCP propagates the darkness clue prompts across all semantic layers of the foundation tracker. Moreover, we design a gated feature aggregation (GFA, in III-D) mechanism to efficiently fuse bottom-up prompts and enable the complementary integration of learned prompts and information from the foundation model. Ultimately, DCPT can effectively inject learned darkness clue prompts into a frozen daytime model with only a small number of prompt learning-related parameters, obtaining superior nighttime tracking performance.
We summarize our major contributions as follows:
@@
Extensive experiments on four nighttime tracking benchmarks validated the effectiveness of DCPT. Qualitative and quantitative results demonstrate the superiority of DCPT as a nighttime tracker for aerial robots. e.g., on DarkTrack2021 [25], DCPT boosts the base tracker by 4.9% success score with 3.0M prompting parameters.
@@ II-A Nighttime UAV Tracking
The emergence of deep-learning technologies and efforts by researchers push visual object tracking to new frontiers[22, 23, 5, 6]. Impressive tracking performance was also achieved on the UAV platforms [24, 4]. However, these trackers, which are mainly designed for daytime scenarios, often suffer severe performance degradation or even fail to work when faced with common but challenging nighttime scenarios. This is because the nighttime circumstances suffer from loss of detailed information, accompanying noise, and low contrast and brightness. Therefore, nighttime UAV tracking has attracted increasing attention. A straightforward manner is to complete nighttime tracking by an "Enhance-then-Track" process. Specifically, researchers design low light enhancers [25, 15, 14] for nighttime scenarios, and use existing daytime UAV trackers to track in the enhanced sequences. Li et al. [16] propose to integrate a low-light image enhancer into a CF-based tracker for robust tracking at night. Similarly, DarkLighter [15] and HighlightNet [14] also design low-light image enhancers to alleviate the influence of extreme illumination and highlight the potential objects, respectively. Although effective, the main drawbacks of such approaches are that the nighttime tracker requires an additional trained light enhancer for preprocessing, incurring extra computational costs, and the separation of enhancer and tracker is not conducive to building an end-to-end trainable nighttime UAV tracking method. Another approach is to adopt domain adaptation to transfer daytime tracker to nighttime scenarios. To obtain the tracking capabilities in the dark, UDAT [17] proposes to align image features from daytime and nighttime domains by the transformer-based bridging layer, in this way, the tracking capabilities on the daytime domain are somewhat transferred to the nighttime domain. Unfortunately, this approach requires more training costs and the lack of high-quality target domain data for tracking also limits its enhancement.
@@ III-A Overview
The overall architecture is shown in Fig. 2. To summarize, DCPT injects the learned darkness clue prompts into the daytime tracker and stimulates its tracking potential in low-light circumstances. The darkness clue prompts are uncovered through end-to-end prompt learning in nighttime data, having the ability to discriminate object tracks in the darkness, which is used to complement the shortcomings of the daytime tracker. These learned darkness clue prompts are propagated from preceding prompts and foundation feature flows . Specifically, the gated feature aggregation (GFA) mechanism is designed for controlling the weight of these different information sources. In general, the daytime tracker has excellent tracking capabilities on generic scenarios while just lacking specific adaptations for nighttime scenarios. Therefore, we only tune parameters that are related to darkness clue prompt generation instead of fine-tuning the entire model. The DCPT framework maximizes the capabilities inherited from daytime tracker trained on large-scale datasets [11, 27, 12], avoiding overfitting on limited nighttime tracking data but gaining nighttime-specific tracking capabilities through darkness clue prompt learning.
III-B Daytime Foundation Tracker
Daytime and nighttime trackers naturally share a number of basic capabilities, including scene understanding and feature matching between targets, hence, previous nighttime tracking methods [16, 14, 15] all chose cutting-edge trackers [28, 4, 23] as their daytime base trackers. Hereby we adopt a streamlined model to serve as a clear baseline, emphasizing the advantages of our darkness clue prompting tracking framework. Specifically, the base tracker is only composed of two parts: a backbone and a prediction head. The backbone consists of stacked vision transformer (ViT) layers and propagating the inputs in a one-stream style like the method in [6]. First, both template and search region are embedded and flattened to 1D tokens with added positional embedding :
@@ V Conclusion
Acknowledgment
The work is supported by the National Natural Science Foundation of China (No.62206040, 62293540, 62293542).